OpenStudy (anonymous):

Given the geometric sequence where a1 = −1 and the common ratio is 7, what is the domain for n?

OpenStudy (zzr0ck3r):

You want \(7=\dfrac{a_2}{a_1}=\dfrac{a_3}{a_2}=\dfrac{a_4}{a_3}\) And we know \(a_1=-1\) so what do we need \(a_2\) to be? \(7=\dfrac{a_2}{-1}\)

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

ok.

OpenStudy (zzr0ck3r):

\(\dfrac{7}{-1}=7\) ?

OpenStudy (anonymous):

yes